Father bought 4 2/4m of chicken wire. How many 3/4m of chicken wires were there?
Answer by Theo(13342)

father bought (4 + 2/4) meters of chicken wire.

how many 3/4 meters of chicken wires were there.

you would want to divide (4 + 2/4) by (3/4)

that's the same as multiplying (4 + 2/4) * 4/3

that is equal to 4 * 4/3 + 2/4 * 4/3.

that is equal to 16/3 + 8/12.

put everything under a common denominator to get:

that is equal to 64/12 + 8/12 = 72/12

72/12 = 6

that says father will get 6 pieces of chicken wire, each measuring 3/4 meters long.

you can confirm this is true by using your calculator to get (4 + 2/4) / (3/4) = 6.

if you want to work with improper fractions, then 4 + 2/4 = 18/4.

18/4 / 3/4 = 18/4 * 4/3 = 18/3 = 6